Mr jamison buys a washing machine that is listed at $800. It is advertised at 20% discount. The sale price is reduced another 10% if bought before the end of this week. He buys the washing machine before the end of the week. Find the total charge including a 6% sales tax and a $40 delivery charge.
What is the correct way to set this problem up?

Answer by rfer(16322)

800*0.80=640
640*0.90=576
576*1.06=610.56
610.56+40=$650.56